OpenStudy (anonymous):

Please help, Is my answer correct 5x + y = 6 5x + 3y = -4 The y-coordinate of the solution to the system shown is _____. Answer: -5 -1 1/2 1

OpenStudy (anonymous):

Subtract the bottom equation from the top equation, to get: y = 3. Substitute into the top equation: 5x-3(3) = 6, 5x = 6+9, 5x = 15, x = 3.

OpenStudy (anonymous):

tht wht i got